I understand how to solve proportions but what I am having trouble with is if I have a question such as what is 25% of 48. Can you help me?

Found 2 solutions by malikul, pro:
Answer by malikul(1)

Yes i can. Since know that 25% is one quater of 100 so that becomes one/four or it can also become .25.

then after you get that
Step1: first put percent in decimal form
step2: multiply by the number you wanna find it by

then it would be
48
x.25
------
12

Answer by pro(1)

%=100 therefore,25%=25/100=0.25
ans = 0.25*48
= 12